# Large Language Models for Wireless Communications: From Adaptation to Autonomy

### v1

@@ Abstract
The emergence of large language models (LLMs) has revolutionized artificial intelligence, offering unprecedented capabilities in reasoning, generalization, and zero-shot learning. These strengths open new frontiers in wireless communications, where increasing complexity and dynamics demand intelligent and adaptive solutions. This article explores the role of LLMs in transforming wireless systems across three key directions: adapting pretrained LLMs for core communication tasks, developing wireless-specific foundation models to balance versatility and efficiency, and enabling agentic LLMs with autonomous reasoning and coordination capabilities. We highlight recent advances, practical case studies, and the unique benefits of LLM-based approaches over traditional methods. Finally, we outline open challenges and research opportunities—including multimodal fusion, collaboration with lightweight models, and self-improving capabilities—charting a path toward intelligent, adaptive, and autonomous wireless networks of the future.
@@ I Introduction
Next-generation wireless systems are characterized by growing complexity and variability. Technologies, such as massive multiple-input multiple-output (MIMO), ultra-dense deployments, and semantic-aware communication, introduce high-dimensional, time-varying environments with stringent latency and reliability constraints. Traditional signal processing and optimization-based methods struggle to handle this scale and non-stationarity. At the same time, existing machine learning models are typically tailored to specific tasks, lack robustness to distribution shifts, and require retraining when system configurations change. LLMs, by contrast, offer general-purpose reasoning, strong transferability, and cross-task knowledge sharing—making them a promising foundation for more flexible and resilient wireless intelligence. However, harnessing LLMs for wireless tasks is non-trivial. A fundamental modality gap exists between the discrete, semantic token space of LLMs and the continuous, high-dimensional numerical data of wireless systems, such as complex-valued channel state information (CSI) or beamforming vectors. Furthermore, LLMs are typically autoregressive and computationally intensive, posing challenges for low-latency inference, particularly in real-time systems.
@@
II Adapting LLMs to Core Wireless Tasks
@@ II-A Physical Layer Prediction Tasks
Figure 1: Network architecture of LLM-based beam prediction [1].
The key challenge in this adaptation lies in the modality mismatch: standard LLMs are designed to operate on discrete token sequences derived from natural language, whereas wireless data spans a range of formats—from continuous, high-dimensional tensors, e.g., CSI, to discrete but domain-specific labels e.g., beam indices. To bridge this gap, recent work has focused on modifying the input and output interfaces of LLMs. As illustrated in Fig. 1, the BP-LLM framework [1] introduces a dual-input strategy: a natural language prompt is used to activate the model’s reasoning capability, while historical beam and angle-of-departure (AoD) sequences are converted into token-like embeddings via a trainable tokenizer. The output is produced not through the default softmax layer used for text generation, but through a task-specific linear head designed to predict beam indices. In contrast, to adopt a different approach for channel prediction, the method in [2] bypasses the language interface altogether by injecting learned CSI embeddings directly into the attention layers of the LLM, effectively treating signal features as first-class model inputs. To preserve the generalization benefits of pretraining while minimizing finetuning costs, these adaptations often rely on parameter-efficient tuning strategies, such as low-rank adaptation or adapters.
Experimental results from the above studies show that adapted LLMs not only achieve competitive performance in full-data regimes but also excel in few-shot and zero-shot settings. Their robustness across varying environments, such as changes in frequency bands or deployment topologies, demonstrates the potential of LLMs to generalize beyond the traditional models, which often suffer from brittle performance under distribution shifts. These findings underscore the promise of LLM-based architectures as flexible and generalizable predictors for core physical layer tasks in wireless systems.
@@
While the adaptation of LLMs to the physical layer requires overcoming substantial modality mismatches, recent advances in tokenization, embedding design, and efficient finetuning have demonstrated their potential to serve as robust, generalizable predictors for a wide range of physical layer tasks. Their ability to transfer across environments and tasks makes them a compelling alternative to traditional learning-based models in next-generation wireless systems.
@@ II-B Resource Allocation
Figure 2: Illustration of personalized semantic extraction via LLM in the semantic communication system for group photo transmission. Currently, Justin (sender) transmits this photo to Jane (receiver), and both of them intends to identify people and background in this photo. In the sender side, the LLM uses their prompt bases to extract intention-aligned semantics (omitting irrelevant details like clothing colors). Furthermore, the LLM transforms sender-centric references to receiver-perspective semantics (’me’ to ’Justin’, ’Jane’ to ’me’). Due to their intention only on people and background, clothing colors in the recover image differ from the original. New users David and Helen can seamlessly join in this system by providing their prompts focused on clothing. The LLM now prioritizes clothes details over background, yielding a recovered image for Helen that preserves original clothing colors but alters the background.
@@
Following input processing, the output heads map the feature representations to specific resource allocation actions, such as transmit power, precoding vectors, and scheduling or spectrum allocation indicators. However, the sequential output process of standard LLMs introduces two critical issues for real-time applications: significant inference latency and the risk of generating invalid outputs that fall outside the defined action space (“hallucinations”). To address this, architectural innovations, such as NetLLM, replace the language-based output head with a task-specific classifier, which maps internal representations directly to the vaild action space [6]. This design enables the model to output a valid action in a single inference step, thereby meeting the strict latency and reliability demands of real-time systems.
@@ II-C Semantic Communication
This paradigm shift enables greater flexibility. For example, in the multi-user generative semantic communication framework [7], the LLM can interpret the task as requiring an object-focused scene layout when prompted with a scene reconstruction task. This interpretation guides the LLM to choose an object detector as the encoder for object detection, and a diffusion model as the decoder to reconstruct the scene. In the large AI model-based multimodal semantic communication system [8], user-specific prompts guide semantic extraction and recovery, supporting personalized semantics based on identity, interests, and intent. As shown in Fig. 2, the LLM adjusts both the encoded content and receiver interpretation using personalized prompt bases, allowing multiple users to perceive the same message differently based on their preference.
@@ III Wireless Foundation Model
The concept of foundation models has reshaped DL in recent years [9], emphasizing general-purpose architectures pretrained on diverse data from a target domain. Unlike massive LLMs, wireless foundation models are not defined by size alone. Rather, their key advantage lies in their ability to encode domain-specific knowledge through pretraining, enabling few-shot, zero-shot, and multi-task capabilities across a variety of communication scenarios, while maintaining fast inference. In what follows, we examine the development of wireless foundation models for two core areas: physical layer prediction and resource management.
@@ III-B Predictive Foundation Models for Resource Management
While Section II discussed LLMs for resource allocation as decision-makers, foundation models serve a different but equally critical role: generating predictive inputs that drive better decisions. Resource management in wireless systems—spanning tasks, such as user scheduling, handover, and power control—depends heavily on accurate forecasting of channel conditions, user mobility, and traffic demand [4]. However, existing predictive models are designed for specific scenarios and struggle to generalize to new environments or support long-horizon predictions.
@@
More importantly, these predictions can be seamlessly integrated into RL frameworks. In model-free RL, the predicted multi-step future information can be embedded into the agent’s state representation, allowing it to optimize for long-term objectives, such as energy efficiency, rather than short-term utility. In model-based RL, predictive trajectories can be used to simulate future states and evaluate the outcome of different actions. For example, Monte Carlo tree search (MCTS) [13] can leverage these predictions to plan multiple steps ahead, which is particularly useful in wireless systems where control overhead and latency are tightly constrained. By coupling predictive foundation models with RL, researchers can develop intelligent and scalable solutions for adaptive resource management in complex and dynamic environments.
@@ III-C Case Study: Zero-Shot Generalization in Multi-Task Predictive Foundation Model
This case study investigates the wireless foundation model in [12] that performs multi-task prediction for three different variables of interest: CSI, AoA, and network traffic. To evaluate its zero-short generalization capability, we directly test the model on an unseen task: predicting user distance, without any additional training. As shown in Fig. 4, the foundation model (“Foundation Model (Zero-shot)”) outperforms a standard full-shot LSTM baseline, trained under the user distance prediction task, in normalized mean-squared error (NMSE), despite never being trained on the distance prediction task. After minimal finetuning (“Foundation Model (Finetuning)”), performance improves further. These results highlight the strong zero-shot learning capabilities of the wireless foundation model.
@@ IV Agentic LLM for Wireless Communication
The preceding sections have shown how pretrained LLMs and wireless foundation models can enhance communication functions, such as semantic compression, channel prediction, and resource management. These advances primarily address well-defined tasks with clear inputs, outputs, and performance goals. However, 6G networks will introduce far greater scale, dynamism, and service heterogeneity, where decision-making becomes open-ended and context-dependent. Static protocols and task-specific models will struggle to adapt to this complexity. To meet these demands, a promising direction is to endow LLMs with agency—the ability to perceive, reason, and act autonomously in dynamic environments.
@@ IV-A Key Capabilities of Agentic LLMs
An agent is generally defined as an entity that perceives its environment, takes autonomous actions toward a goal, and improves its behavior over time through learning. Traditional approaches include rule-based systems and RL. While rule-based agents are interpretable, they lack flexibility and scalability. RL agents, though adaptive, often start with no prior knowledge, require a large amount of exploration data, and tend to overfit to specific training environments—limiting their generalization capacity in dynamic wireless networks.
Agentic LLMs overcome these limitations by combining the reasoning ability of symbolic systems, the adaptability of RL, and the prior knowledge in LLMs. Three core capabilities underpin their effectiveness in wireless communication:
  • Reasoning and Planning: Agentic LLMs can parse complex task prompts, decompose them into subgoals, and plan corresponding actions using natural language. Unlike traditional RL agents that rely on repeated environment interaction to learn policies, LLM-based agents can formulate strategies in zero-shot settings by leveraging their pretrained knowledge and reasoning abilities.
  • Memory and Reflection: Memory plays a central role in enabling long-term adaptation. Agentic LLMs utilize a dual memory system: short-term memory (the context for ongoing decisions) and long-term memory (accumulated knowledge, past interactions, and retrieved facts, e.g., via retrieval-augmented generation [14]). These models can reflect on past actions, identify errors or suboptimal decisions, and update memory accordingly.
  • Tool Use: Tool use equips agentic LLMs with the ability to interact with external environments through standardized interfaces and protocols. These interfaces enable the invocation of external tools, including channel simulators, resource allocation algorithms, and other APIs, extending the agent’s perceptual and actuation capabilities.
@@ IV-B Applications in Wireless Systems
Leveraging the key capabilities of agentic LLMs, recent work has demonstrated their potential in practical wireless scenarios. The WirelessAgent prototype [15] showcases how an agentic LLM can infer user intent, optimize in-slice beamforming, and dynamically manage network slicing to meet quality of service requirements. This illustrates a pathway toward self-organizing networks, where an agentic LLM functions as a centralized “brain” with full network awareness, memory of traffic patterns, and the ability to predict future service needs.
In this paradigm, traditionally fragmented tasks, spanning protocol layers, are unified under a single decision-making framework. The agent interprets a natural language prompt that specifies cross-layer constraints, including latency budgets, energy limits, and service priorities. It then leverages tools, such as channel simulators, to evaluate potential configurations. Through a process of reflection, the agent iteratively refines its strategy. This iterative process enables the agent to converge on high-quality solutions without the extensive environment interaction required by conventional RL approaches. However, as wireless networks scale, the centralized “brain” architecture faces bottlenecks in latency, computation, and fault tolerance. To address this, collaborative designs involving multiple agentic LLMs have been proposed, typically following two structural paradigms.
In vertical architectures, a leader agent aggregates local reports, performs global reasoning, and dispatches control commands. This top-down model supports fast convergence and clear coordination but introduces a single point of failure and limited scalability tied to the lead agent’s capacity.
In contrast, horizontal architectures distribute autonomy: all agents share a communication interface, broadcast observations, negotiate sub-tasks, and invoke tools independently. This peer-to-peer structure improves fault tolerance and scalability—suitable for decentralized use cases like federated learning and interference coordination—but may suffer from excessive communication overhead if not optimized through message compression or clustering strategies. Overall, balancing and integrating these architectural patterns will be essential to developing resilient, scalable multi-agent systems capable of powering the next generation of intelligent wireless networks.
@@ IV-C Case Study: Multi-AP Coordination with Agentic LLMs
This case study illustrates a vertical multi-agent framework using agentic LLMs for coordinated transmission in the Wi-Fi overlapping basic service set, overlapping BSS or OBSS, environments. Each access point (AP) is modeled as an autonomous LLM agent capable of natural language communication, reasoning, and decision-making, thereby enabling decentralized cooperation in dense wireless deployments.
Recent updates in Wi-Fi 8 introduce standardized support for multi-AP coordination, including mechanisms, such as coordinated time division multiple access (Co-TDMA) and coordinated spatial reuse (Co-SR). These mechanisms allow APs to share critical information, such as CSI and link configurations, to mitigate inter-BSS interference and improve spectrum utilization. In our agentic framework, these coordination procedures are facilitated through communication among AP agents, allowing them to negotiate and learn adaptive channel access strategies in real time.
Fig. 5 outlines the agentic coordination protocol. APs first contend for channel access using the standard carrier-sense multiple access with collision avoidance (CSMA/CA) procedure. Once a coordination group is formed, a designated sharing AP leads the communication round, occupying the channel for a transmission opportunity (TXOP). During this TXOP, the sharing AP interacts with the shared APs through agentic dialogue to collaboratively determine slot assignments and transmission strategies. Each AP-agent leverages reasoning, memory, and reflection modules to update its local policy, adapting to dynamic environmental conditions and past coordination outcomes.
@@
Simulation results in Fig. 6 validate the effectiveness of the proposed protocol. The agentic LLMs progressively learn interference-aware and topology-adaptive strategies, achieving efficient channel utilization across varying OBSS configurations. Notably, the agents exhibit strong environmental adaptability, a capability enabled by their natural language-based generalization and reasoning. Moreover, the framework imposes no rigid heuristics, allowing agents to self-organize and negotiate actions based on shared intent and real-time conditions. This flexibility suggests a promising path toward scalable multi-agent coordination in future wireless systems, especially in scenarios where traditional rule-based protocols may fail to adapt quickly to environmental dynamics.

### v2

@@ Abstract
The emergence of large language models (LLMs) has revolutionized artificial intelligence, offering unprecedented capabilities in reasoning, generalization, and zero-shot learning. These strengths open new frontiers in wireless communications, where increasing complexity and dynamics demand intelligent and adaptive solutions. This article explores the role of LLMs in transforming wireless systems across three key directions: adapting pretrained LLMs for communication tasks, developing wireless-specific foundation models to balance versatility and efficiency, and enabling agentic LLMs with autonomous reasoning and coordination capabilities. We highlight recent advances, practical case studies, and the unique benefits of LLM-based approaches over traditional methods. Finally, we outline open challenges and research opportunities—including multimodal fusion, collaboration with lightweight models, and self-improving capabilities—charting a path toward intelligent, adaptive, and autonomous wireless networks.
@@ I Introduction
Next-generation wireless systems are characterized by growing complexity and variability. Technologies, such as massive multiple-input multiple-output (MIMO), ultra-dense deployments, and semantic-aware communication, introduce high-dimensional, time-varying environments with stringent latency and reliability constraints. Firstly, these specialized models exhibit poor generalization, often failing to adapt when deployed in new wireless environments with different channel statistics. Secondly, their limited capacity imposes a performance ceiling on increasingly complex tasks where capturing intricate data patterns is essential. Thirdly, their inherent inflexibility makes them prone to overfitting on a single objective, rendering them unsuitable for dynamic, multi-faceted problems. LLMs, by contrast, offer general-purpose reasoning, strong transferability, and cross-task knowledge sharing—making them a promising foundation for more flexible and resilient wireless intelligence. However, harnessing LLMs for wireless tasks is non-trivial. There exists fundamental modality gap between the discrete, semantic token space of LLMs and the continuous, high-dimensional numerical data of wireless systems, such as complex-valued channel state information (CSI) or beamforming vectors. Furthermore, LLMs are typically autoregressive and computationally intensive, posing challenges for low-latency inference in real-time systems.
@@
II Adapting LLMs to Wireless Tasks
@@ II-A Physical Layer Prediction Tasks
Figure 1: Network architecture of LLM-based beam prediction.
The key challenge in this adaptation lies in the modality mismatch: standard LLMs are designed to operate on discrete token sequences derived from natural language, whereas wireless data spans a range of formats—from continuous, high-dimensional tensors, e.g., CSI, to discrete but domain-specific labels e.g., beam indices. To bridge this gap, recent work has focused on modifying the input and output interfaces of LLMs, leveraging their high-dimensional hidden states to create rich token embeddings that can represent complex signal features with minimal loss of fidelity. As illustrated in Fig. 1, the beam prediction framework based on LLMs (BP-LLM) [1] introduces a dual-input strategy: a natural language prompt is used to activate the model’s reasoning capability, while historical beam and angle-of-departure (AoD) sequences are converted into token-like embeddings via a trainable tokenizer. The output is produced through a task-specific linear head designed to predict beam indices. In contrast, to adopt a different approach for channel prediction, the method in [2] bypasses the language interface altogether by injecting learned CSI embeddings directly into the attention layers of the LLM, effectively treating signal features as first-class model inputs. To preserve the generalization benefits of pretraining while minimizing finetuning costs, these adaptations often rely on parameter-efficient tuning strategies, such as low-rank adaptation or adapters.
Experimental results from the above studies show that adapted LLMs not only achieve competitive performance in full-data regimes but also excel in few-shot and zero-shot settings. Their robustness across varying environments, such as changes in frequency bands or deployment topologies, demonstrates the potential of LLMs to generalize beyond traditional models, which often suffer from brittle performance under distribution shifts. These findings underscore the promise of LLM-based architectures as flexible and generalizable predictors for core physical layer tasks in wireless systems.
@@
While the adaptation of LLMs to the physical layer requires overcoming substantial modality mismatches, recent advances in tokenization, embedding design, and efficient finetuning have demonstrated their potential to serve as robust, generalizable predictors for a wide range of physical layer tasks. The primary hurdle to their deployment in real-time systems, inference latency, can be mitigated by promising trends in lightweight architectures, dedicated hardware acceleration, and domain-specific foundation models. This progress, combined with their ability to transfer across environments and tasks, makes them an alternative to traditional learning-based models in future 6G networks.
@@ II-B Resource Allocation
Figure 2: Illustration of personalized semantic extraction via LLM in the semantic communication system for group photo transmission. Currently, Justin (sender) transmits this photo to Jane (receiver), and both of them intend to identify people and background in this photo. In the sender side, the LLM uses their prompt bases to extract intention-aligned semantics (omitting irrelevant details like clothing colors). Furthermore, the LLM transforms sender-centric references to receiver-perspective semantics (“me” to “Justin”, “Jane” to “me”). Due to their intention only on people and background, clothing colors in the recovered image differ from the original. New users David and Helen can seamlessly join in this system by providing their prompts focused on clothing. The LLM now prioritizes clothes details over background, yielding a recovered image for Helen that preserves original clothing colors but alters the background.
@@
Following input processing, the output heads map the feature representations to specific resource allocation actions, such as transmit power, precoding vectors, and scheduling indicators. However, the sequential output process of LLMs introduces two critical issues for real-time applications: significant inference latency and the risk of generating invalid outputs that fall outside the defined action space (“hallucinations”). To address this, architectural innovations, such as NetLLM, replace the language-based output head with a task-specific classifier, which maps internal representations directly to the vaild action space [6]. This design enables the model to output a valid action in a single inference step, thereby meeting the strict latency and reliability demands of real-time systems.
@@ II-C Semantic Communication
This paradigm shift enables greater flexibility. For example, in the multi-user generative semantic communication framework [7], the LLM can interpret the task as requiring an object-focused scene layout when prompted with a scene reconstruction task. In the large AI model-based multimodal semantic communication system [8], user-specific prompts guide semantic extraction and recovery, supporting personalized semantics based on identity, interests, and intent. As shown in Fig. 2, the LLM adjusts both the encoded content and receiver interpretation using personalized prompt bases, allowing multiple users to perceive the same message differently based on their preference.
@@ III Wireless Foundation Model
The concept of foundation models has reshaped DL in recent years [9], emphasizing general-purpose architectures pretrained on diverse data from a target domain. Unlike LLMs, wireless foundation models are not defined by size alone. Rather, their key advantage lies in their ability to encode domain-specific knowledge through pretraining, enabling few-shot, zero-shot, and multi-task capabilities across different communication scenarios, while maintaining fast inference. In what follows, we examine the development of wireless foundation models for two core areas: physical layer prediction and resource management.
@@ III-B Predictive Foundation Models for Resource Management
While Section II discussed LLMs for resource allocation as decision-makers, foundation models serve a different but equally critical role: generating predictive inputs that drive better decisions. Resource management in wireless systems—including user scheduling, handovers, and power control—depends heavily on accurate forecasting of channel conditions, user mobility, and traffic demand [4]. However, existing predictive models are often scenario-specific, exhibiting limited generalization to new environments and struggling with long-horizon predictions.
@@
More importantly, these predictions can be seamlessly integrated into RL frameworks. In model-free RL, the predicted multi-step future information can be embedded into the agent’s state representation, enabling optimization for long-term objectives like energy efficiency over short-term utility. In model-based RL, these predictive trajectories allow for simulating future states to evaluate different actions. For example, Monte Carlo tree search (MCTS) [13] can leverage these predictions to plan multiple steps ahead, which is particularly useful in wireless systems where control overhead and latency are tightly constrained. This synergy between predictive foundation models and RL enables the development of intelligent and scalable solutions for adaptive resource management in complex, dynamic environments.
@@ III-C Case Study: Zero-Shot Generalization in Multi-Task Predictive Foundation Model
This case study investigates the wireless foundation model in [12] that performs multi-task prediction for three different variables of interest: CSI, AoA, and network traffic. To evaluate its zero-shot generalization capability, we directly test the model on an unseen task: predicting user distance, without any additional training. As shown in Fig. 4, the foundation model (“Foundation Model (Zero-shot)”) outperforms the standard full-shot LSTM and Transformer baselines, trained under the user distance prediction task, in normalized mean-squared error (NMSE), despite never being trained on the distance prediction task. This superior zero-shot performance stems from the model’s ability to capitalize on the inherent structural similarities among these time-series prediction tasks. Extensive pre-training yields a generalized understanding of temporal dynamics, which proves directly applicable to the unseen distance prediction problem. After minimal finetuning (“Foundation Model (Finetuning)”), performance improves further. These results highlight the strong zero-shot learning capabilities of the wireless foundation model.
@@ IV Agentic LLM for Wireless Communication
The preceding sections have shown how pretrained LLMs and wireless foundation models can enhance communication functions, such as semantic compression, channel prediction, and resource management. These advances primarily address well-defined tasks with clear inputs, outputs, and performance goals. However, to manage the immense complexity of future 6G networks, a paradigm shift is required from creating task-specific tools to engineering autonomous systems. This marks the final stage in the evolution from adaptation to autonomy: endowing LLMs with agency—the ability to perceive, reason, and act autonomously in dynamic environments. While the foundational concepts of agentic AI are rooted in general computer science, our focus here is their integration into wireless systems to achieve holistic network self-organization.
@@ IV-A Key Capabilities of Agentic LLMs
Unlike traditional approaches such as rule-based systems or pure RL agents, which often lack flexibility or require extensive environment-specific training, agentic LLMs overcome these limitations by combining the reasoning ability of symbolic systems, the adaptability of RL, and the vast prior knowledge embedded within LLMs.
Crucially, their decision-making policies are often not learned from pre-existing, static datasets. Instead, agentic LLMs can operate through two primary modes: i) in-context learning, where strategies are guided by detailed prompts, and ii) online interaction with the environment (or a simulator). In the latter mode, agents generate their own streams of experience and adapt their policies via reflection, a paradigm recently explored for intelligent network management [14]. The effectiveness of this approach in wireless communication is underpinned by three core capabilities:
  • Reasoning and Planning: The ability to parse high-level goals, decompose them into a sequence of executable subgoals, and formulate a multi-step action plan. This core capability enables them to devise strategies in zero-shot settings by leveraging their pretrained knowledge.
  • Memory and Reflection: The capacity to learn from past interactions by utilizing a dual memory system: short-term memory for immediate context and long-term memory for accumulated knowledge [15]. This allows agents to reflect on past actions, identify errors, and continually refine their strategies over time.
  • Tool Use: The functionality to interact with external environments through standardized interfaces. This enables the invocation of specialized tools—such as channel simulators, resource allocation algorithms, and other APIs—thereby extending the agent’s perceptual and actuation capabilities beyond its internal knowledge.
@@ IV-B Applications in Wireless Systems
Recent work has demonstrated the potential of agentic LLMs in practical wireless applications. The WirelessAgent prototype [14] showcases how an agentic LLM can infer user intent, optimize in-slice beamforming, and dynamically manage network slicing to meet quality of service requirements. This illustrates a pathway toward self-organizing networks, where agentic LLMs operate as a centralized controller with comprehensive network awareness, memory of traffic patterns, and predictive capabilities for future service demands.
This centralized model unifies traditionally fragmented, cross-layer tasks into a single decision-making framework. The agent interprets natural language prompts specifying cross-layer constraints, including latency budgets, energy limits, and service priorities. It then leverages tools, such as channel simulators, to evaluate potential configurations. Through reflection, the agent iteratively refines its strategy, converging on high-quality solutions without the extensive environment interaction demanded by conventional RL approaches. However, as wireless networks scale, this centralized architecture encounters critical bottlenecks in latency, computation, and fault tolerance. These limitations necessitate collaborative multi-agent designs, which typically follow two structural paradigms.
In vertical architectures, a leader agent aggregates local reports, performs global reasoning, and dispatches control commands. This top-down model supports fast convergence and clear coordination but introduces a single point of failure and limited scalability tied to the lead agent’s capacity. In contrast, a horizontal architecture distributes autonomy by allowing all agents to share a communication interface, broadcast observations, negotiate sub-tasks, and invoke tools independently. This peer-to-peer structure enhances fault tolerance and scalability, making it suitable for decentralized tasks like federated learning and interference coordination. However, it can suffer from excessive communication overhead unless optimized with strategies such as message compression or clustering. Integrating these architectural patterns is crucial for developing resilient and scalable multi-agent systems to power next-generation intelligent wireless networks.
@@ IV-C Case Study: Multi-AP Coordination with Agentic LLMs
This case study illustrates a decentralized vertical multi-agent framework using agentic LLMs for coordinated transmission in the Wi-Fi overlapping basic service set (OBSS) environments. Each access point (AP) is modeled as an autonomous LLM agent capable of natural language communication, reasoning, and decision-making, thereby enabling adaptive cooperation in dense wireless deployments.
Recent updates in Wi-Fi 8 introduce standardized support for multi-AP coordination, including mechanisms, such as coordinated time division multiple access (Co-TDMA) and coordinated spatial reuse (Co-SR). These mechanisms allow APs to share critical information to mitigate interference between basic service sets (BSSs). In our agentic framework, these coordination procedures are facilitated through communication among AP agents, allowing them to negotiate and learn adaptive channel access strategies.
Figure 5 outlines the agentic workflow for sharing AP and the coordination protocol for multi-AP cooperation. Internally, each agent continuously refines its strategy through a cognitive loop of evaluation and reflection. Leveraging memory of past outcomes to evaluate performance, the agent reflects on the outcomes of its actions (e.g., transmission collisions) to proactively refine its strategy. External tools ground this internal process in the physical world. Externally, this reflective capability enables a sophisticated coordination protocol where agents engage in multi-round negotiation and feedback, allowing them to collectively adapt to the environment rather than adhering to rigid, pre-defined rules.
@@
Simulation results in Fig. 6 validate the effectiveness of the proposed protocol. In the Co-TDMA scenario, the agents initially experience several collisions. However, by leveraging the reflective and negotiation processes detailed in Fig. 5, they rapidly converge on a stable, collision-free transmission schedule after a few rounds. In the Co-SR scenario, the agents progressively explore more aggressive strategies to maximize throughput. This ability to autonomously discover and adapt coordination strategies, without reliance on pre-programmed heuristics, underscores the potential of agentic LLMs to manage dynamic and complex wireless environments.